OpenStudy (anonymous):

Solve the following system of inequalities by graphing. 5x - 4y ≥ -20 3x - 2y ≤ -8 y ≥ -3

OpenStudy (anonymous):

you hve graphs? y <5/4x +5 y> 3/2x +4 with the equl underneath 1. graph the y-interceept 2. from y- intercept use rise over run or slope. 3.connect since you have equal to sign solid line 4. shade above or below the line easy way to know your right p[lug in origin.

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

Solve the following system of inequalities by graphing. 5x - 4y ≥ -20 3x - 2y ≤ -8 y ≥ -3 11. Solve the following system of inequalities by graphing. -9x + 6y ≥ 18 -x + 3y ≥ -12 y ≤ 4 12. Solve the following system of inequalities by graphing. 3x + 12y ≥ -108 -7x + 6y ≤ 48 y ≤ -4 13. Solve the following system of inequalities by graphing. 6x + 10y ≤ -40 x + y ≤ -8 y ≥ -1 14. Solve the following system of inequalities by graphing. -4x + 16y ≥ 128 -3x + 8y ≤ 80 y ≤ 12 15. Solve the following system of inequalities by graphing. 3x + 9y < -18 -x + 3y ≤ 12 y < 0 16. Solve the following system of inequalities by graphing. 5x + 2y ≥ -6 -5x + 2y < 14 y < 6 17. Solve the following system of inequalities by graphing. -6x - 2y < -16 2x - y > 2 y ≥ -4 Part 3 Select the best answer from the choices provided. (Each question is worth one point) 18. Lee and Pat have $40. Lee has three times the money that Pat has. Which of the following is a system of equations that describes this situation? 4L + P = 40 3L = P L + 4P = 40 L = 3P L + P = 40 L = 3P L + P = 40 L = -3P 19. A small toy company makes two types of toys: dolls and toy cars. The expense to make a toy car is $5. The expense to make a doll is $7. The company spends $780 a day making toys. They also make 20 toy cars fewer than twice the number of dolls they make. Construct a system of equations that describes this situation. 7c + 5d = 780 c + 20 = -2d 7c + 5d = 780 c - 20 = -2d 5c + 7d = 780 c + 20 = 2d 5c - 7d = 780 c – 20 = 2d 20. Susan and Joyce have $45. Susan has ten dollars more than four times the money that Joyce has. Construct a system of equations that describes this situation. S + J = 45 S + 10 = -4J S + J = 45 S - 10 = 4J S + J = 45 S + 10 = 4J S + J = 10 S + 45 = 4J 21. Tim has $200 more than Ron. Tim has 50 dollars more than twice the money that Ron has. Construct a system of equations that describes this situation. T + R = 200 T + 2R = 50 T - R = 200 T + 2R = 50 T - R = 200 T – 2R = 50 R - T = 200 R – 2T = 50 Part 4 Type the answer to the question in the textbox below each item. (Each question is worth two points) 22. Solve the following systems of equations for both x and y. -5x + 3y = -40 9x - 4y = 72 23. Solve the following systems of equations for both x and y. -10x + 9y = 73 -5x - 2y = 56 24. Solve the following systems of equations for both x and y. -x - 2y = -17 2x + y = 16 Save My Work So I Can Continue Later Submit My Answers

OpenStudy (anonymous):

@poetrylover

OpenStudy (anonymous):

okay give me some time okay?

OpenStudy (anonymous):

follow these steps 1. isolate y to get it in y > or < mx +b form okay. 2. graph y-intercept and from there do slope rise over run 3. depending on the sign greater than or equal to as well as less than or equal to you will have a solid line. greater than or less than you have a dash line. 4. to shade plug in (0,0) for x and y until you make the equation true.

OpenStudy (anonymous):

i'll show you an ex